A child leaves her book bag on a slide. The bag, w
Answer #1

a) a = 2L/t2 = 2 x 3.7/(1.73)2 = 2.4725 m/s2

b) friction force = mgsin31 - ma = 3.3 x 9.8 x sin31 - 3.3 x 2.4725 = 8.497 N

c) u = F/mgcos31 = 8.497/(3.3 x 9.8 x cos31) = 0.3065

d) speed = at = 2.4725 x 1.73 = 4.2775 m/s